# Bayesian Perspective for Orientation Determination in Cryo-EM with Application to Structural Heterogeneity Analysis

### v2

@@ 2.3 The MLE, MAP and MMSE orientation estimators
In the following presentation, we focus on independent and identically distributed Gaussian noise with $\Sigma=\sigma^{2}I_{d\times d}$. These noise statistics are commonly employed in modern software tools kimanius2021new . While the proposed framework can incorporate more advanced noise models, we restrict attention to this case for clarity of presentation and to simplify the theoretical analysis.
@@
This optimization problem is known as the Orthogonal Procrustes problem in the linear algebra literature. It can be solved efficiently using the singular value decomposition (SVD) of $A$ gower2004procrustes .
Thus, applying the Orthogonal Procrustes procedure to $\hat{g}_{\mathsf{relax}}$ projects it back onto $\mathsf{SO}(3)$, yielding a valid rotation matrix. The following proposition establishes that this valid rotation matrix coincides with the MMSE estimator $\hat{g}_{\mathsf{MMSE}}$, thereby completing the procedure for obtaining the MMSE estimate. A detailed proof is provided in Appendix C.
@@
We conclude the section by showing the next proposition. It shows that the MMSE and MAP estimators converge in the high SNR regime (i.e., $\sigma\to 0$), as demonstrated empirically in Figures 3, 4, and 5. This implies that the MMSE estimator’s superior statistical properties are most advantageous in low SNR conditions, which are common in structural biology applications like cryo-EM and cryo-ET bendory2020single . In these low SNR environments, the MMSE estimator consistently outperforms its MAP counterpart.
@@
The proof of the proposition is provided in Appendix D. The proof relies on the existing result, (robert1999monte, , Theorem 5.10), and is valid not only to rotations but also to other group operators, such as translations.
@@ 3 Numerical methods for MMSE orientation estimation
Since the posterior distribution of $g\in\mathsf{SO}(3)$ in (2.6) is continuous, a numerical discretization of the rotation group is required. In this work, the expectation is approximated using a quadrature rule over $\mathsf{SO}(3)$. Let $\{(g^{(\ell)},w_{\ell})\}_{\ell=0}^{L-1}$ denote a numerical quadrature on $\mathsf{SO}(3)$, constructed as a product of a spherical quadrature on $\mathsf{S}^{2}$ and a circular rule over the Euler angles graf2009sampling ; graf2012unified . The implementation employed here follows the publicly available MATLAB code described in hoskins2024subspace .
@@
Both estimators require evaluating $\|y-x_{\ell}\|$ for all candidate rotations, which naively costs $O(d)$ per candidate and thus $O(Ld)$ overall. This estimate, however, overlooks the structure of $\mathsf{SO}(3)$. Any 3D rotation can be decomposed into a viewing direction (two degrees of freedom, discretized with about $L^{2/3}$ samples) and an in-plane rotation about that direction (one degree of freedom, discretized with $L^{1/3}$ samples). By exploiting FFT-based methods, all $L^{1/3}$ in-plane rotations can be evaluated simultaneously for each viewing direction. Consequently, the effective complexity becomes $O(L^{2/3}\,d\,\log d)$, which improves over the naive bound by a factor of $L^{1/3}$ kostelec2008ffts ; kileel2024fast . This reduction is substantial in practice, since $L$ often reaches tens or even hundreds of thousands in high-resolution cryo-EM.
@@
One key advantage of the Bayesian framework is its flexibility in incorporating different prior distributions for the rotations. In practical cryo-EM applications, the distribution of particle orientations is often non-uniform due to preferred particle orientations or sample preparation artifacts tan2017addressing ; lyumkis2019challenges . If the rotation distribution can be well estimated in an early stage, the information of the rotation distribution can then be integrated into the Bayes estimator to improve rotation estimation.
To illustrate rotation estimation under different prior distributions on $\mathsf{SO}(3)$, we replace the uniform distribution on $\mathsf{SO}(3)$ with an isotropic Gaussian (IG) distribution on $\mathsf{SO}(3)$, denoted by $g\sim\mathcal{IG}_{\mathsf{SO}(3)}(\eta)$, parameterized by a scalar variance $\eta^{2}$. This distribution is frequently used in machine learning probabilistic models on $\mathsf{SO}(3)$ corso2022diffdock ; leach2022denoising ; jagvaral2024unified . It is worth noting that the IG distribution serves as a typical example, and similar phenomena as observed here extend to other non-uniform distributions on $\mathsf{SO}(3)$ as well.
The IG distribution $\mathcal{I}\mathcal{G}_{\mathsf{SO}(3)}(\eta)$ can be represented in an axis-angle form, with uniformly sampled axes of rotation and a rotation angle $\omega\in[0,\pi]$. The scalar variance $\eta^{2}$ controls the distribution of the rotation angle $\omega$: as $\eta\to\infty$, the $\mathsf{SO}(3)$ distribution approaches uniformity, whereas as $\eta\to 0$, $\omega$ becomes increasingly concentrated around 0, i.e., the rotation angle around the rotation axis is small. We apply the inverse sampling method to obtain i.i.d. samples from the IG distribution. Further details on this distribution are provided in Appendix E.
@@
Figure 2: Impact of incorporating the prior rotation distribution on estimation accuracy. Simulations are performed for the cryo-ET model (1.2), excluding the projection step. The true rotation distribution is modeled as an isotropic Gaussian $\mathcal{IG}_{\mathsf{SO}(3)}(\eta=0.1)$. Estimation performance is measured using the geodesic distance defined in (2.8). Here, $g$ denotes the true rotation, $\hat{g}_{\mathrm{MLE}}$ is the maximum-likelihood estimator from (3.4), $\hat{g}_{\mathrm{MAP}}$ is the maximum a posteriori estimator from (2.10), and $\hat{g}_{\mathrm{MMSE}}$ denotes the Bayesian minimum mean square error estimator from (3.1). The MMSE estimators are computed assuming isotropic Gaussian priors on $\mathsf{SO}(3)$ with different concentration parameters $\eta\in\{0.5,0.1\}$ (see Appendix E). As $\eta$ decreases, the prior becomes more concentrated and closer to the true underlying distribution, leading to improved accuracy of both the MAP and MMSE estimators. Each data point in the plot is averaged over 3000 Monte Carlo trials using a rotation grid of size $L=2976$. Below the curve plot, we show visual reconstructions of denoised 3D volumes generated using different rotational estimators. The second row (right panel) corresponds to the MMSE estimator that employs the true prior $\mathcal{IG}_{\mathsf{SO}(3)}(\eta=0.1)$. Incorporating this prior clearly enhances rotational accuracy compared to the MLE baseline.
@@ 4 Bayesian orientation estimation as part of the volume reconstruction problem
Structure reconstruction typically follows two primary approaches: hard-assignment or soft-assignment methods, which are generally implemented through iterative refinement. In the hard-assignment approach, each observation is assigned a single orientation based on the highest correlation, and the 3D structure is then reconstructed given the rotations. In contrast, the soft-assignment method assigns probabilities across all possible orientations for each observation, enabling the 3D structure to be recovered as a weighted average of the observations, with weights determined by these probabilities. The iterative application of the soft-assignment procedure aligns with the EM algorithm dempster1977maximum ; sigworth2010introduction , which serves as the core computational method in modern cryo-EM scheres2012relion ; punjani2017cryosparc .
@@ 4.1 Connection to the EM algorithm in volume reconstruction without projection
Unlike the previous orientation estimation model (2.1), we consider the following simplified model for volume reconstruction (see also singer2018mathematics ; bandeira2023estimation ; fan2024maximum ). We observe $M$ i.i.d samples taking the form
@@
where $V$ is 3D volume structure of interest, $\{g_{i}\}_{i=0}^{M-1}\in\mathsf{SO}(3)$ are unknown latent variables following i.i.d. uniform distribution, satisfying $g_{i}\circ V(x)\equiv V(g_{i}^{-1}x)$, and $\{\varepsilon_{i}\}_{i=0}^{M-1}$ are i.i.d isotropic Gaussian noise with variance $\sigma^{2}$. For the case where $V$ represents a 2D image, $\{g_{i}\}_{i=0}^{M-1}\in\mathsf{SO}(2)$ corresponds to in-plane rotations, and the model is used for image recovery (see more in Section 4.2ma2019heterogeneous . The goal is to recover $V$ from the observations $\{y_{i}\}_{i=0}^{M-1}$, treating the rotations $\{g_{i}\}_{i=0}^{M-1}$ as latent variables.
@@
    $\hat{V}^{(t+1)}=$ $\underset{V}{\operatorname*{arg\,min}}\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\|y_{i}-\left(g\circ V\right)\right\|^{2}\mathrm{d}g$
    $=$ $\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)(g^{-1}\circ y_{i})\mathrm{d}g.$ (4.3)
Proposition 4.1.
Let $\hat{V}^{(t+1)}$ be the $(t+1)$-th volume estimator in the EM algorithm as described in Algorithm 2. Then, the M-step update has the form
$\hat{V}^{(t+1)}=\frac{1}{M}\sum_{i=0}^{M-1}\hat{\mathfrak{g}}_{\mathsf{MMSE},i,t}\circ y_{i},$ (4.4)
where $\hat{\mathfrak{g}}_{\mathsf{MMSE},i,t}$ denotes the MMSE rotation operator defined as the posterier mean of $g^{-1}$ with respect to the uniform prior, given observation $y_{i}$ and underlying structure $\hat{V}^{(t)}$, using the same posterior weight as in (2.11):
$\hat{\mathfrak{g}}_{\mathsf{MMSE},i,t}:=$ $\mathbb{E}[g^{-1}|y_{i},\hat{V}^{(t)}]$
$\mathrel{\text{``=''}}$ $\int_{g\in\mathsf{SO}(3)}\ p_{i}^{(t)}(g)g^{-1}\mathrm{d}g,$ (4.5)
where the quoted equality emphasizes that the expectation (or integral) is interpreted as acting on the operator, in contrast to the previous $3\times 3$ posterior matrix mean in (2.11).
The proof of this proposition is presented in Appendix F. In words, the proposition shows that the update rule for the volume structure estimation at iteration $t+1$, given the volume $\hat{V}^{(t)}$, is equivalent to aligning each observation $\left\{y_{i}\right\}_{i=0}^{M-1}$ using the associated MMSE operator $\hat{\mathfrak{g}}_{\text{MMSE},i,t}$, computed based on the reference volume $\hat{V}^{(t)}$, and then averaging the aligned observations (4.4). Thus, the MMSE rotation operator is a key ingredient in the EM algorithm for volume reconstruction.
@@
The MMSE rotation “operator” introduced above is closely related to the MMSE rotation “estimator”, namely the orthogonal Procrustes projection of the matrix-based posterior mean $\hat{g}_{\mathsf{relax}}$ discussed in (2.12). In the matrix formulation, the expectation produces a $3\times 3$ posterior matrix mean. In contrast, in the operator formulation, the expectation acts directly on the data $y_{i}$, effectively performing a posterior-weighted average over $g^{-1}$. While the two formulations differ in representation, the underlying principle is the same: both correspond to a posterior-weighted aggregation of rotations according to the posterior distribution.
Moreover, if the volume is expressed in a suitable basis, such as the spherical harmonics basis, the operator has a corresponding matrix representation (e.g., a block-diagonal matrix of rotation matrices). In this case, the MMSE operator can also be written as a posterior mean of such block-diagonal rotation matrix.
Despite the subtle differences arising from rotation estimation versus structure reconstruction, the previously introduced MMSE rotation estimator (2.11) can still serve as a practical tool for assessing the performance of soft assignments. Concretely, we take the $3\times 3$ posterior rotation matrix mean as in (2.11),
$\hat{g}_{\mathsf{relax},i,t}:=\int_{g\in\mathsf{SO}(3)}\ p_{i}^{(t)}(g)g\mathrm{d}g,$
perform an Orthogonal Procrustes projection to enforce a valid rotation, and then take the inverse to replace $\hat{\mathfrak{g}}_{\mathsf{MMSE},i,t}$. In this way, the MMSE rotation estimator provides a concrete approximation to the soft assignment in structure reconstruction. In numerical experiments, it is compared with the hard assignment, demonstrating superior performance relative to the maximum-likelihood estimator, as discussed below.

### v3

@@ 2.3 The MLE, MAP and MMSE orientation estimators
In the following presentation, we focus on independent and identically distributed Gaussian noise with $\Sigma=\sigma^{2}I_{d\times d}$. These noise statistics are commonly employed in modern software tools Kimanius et al. (2021). While the proposed framework can incorporate more advanced noise models, we restrict attention to this case for clarity of presentation and to simplify the theoretical analysis.
@@
This optimization problem is known as the Orthogonal Procrustes problem in the linear algebra literature. It can be solved efficiently using the singular value decomposition (SVD) of $A$ Gower and Dijksterhuis (2004). Concretely, given a matrix $A\in\mathbb{R}^{3\times 3}$ (in our case $A=\hat{g}_{\mathrm{relax}}$), compute its singular value decomposition
$A=U\Sigma V^{\top},\qquad U,V\in O(3),\ \Sigma\succeq 0.$ (2.15)
The closest rotation matrix in Frobenius norm is then obtained by
$\Omega=U\,\mathrm{diag}\!\bigl(1,\,1,\,\det(UV^{\top})\bigr)\,V^{\top}\ \in\mathsf{SO}(3).$ (2.16)
Equivalently, if $\det(UV^{\top})=-1$, one flips the sign of the last column of $U$ (or $V$) before forming $UV^{\top}$, ensuring $\det(\Omega)=1$ and yielding the Procrustes solution. Thus, applying the Orthogonal Procrustes procedure to $\hat{g}_{\mathsf{relax}}$ projects it back onto $\mathsf{SO}(3)$, yielding a valid rotation matrix. The following proposition establishes that this valid rotation matrix coincides with the MMSE estimator $\hat{g}_{\mathsf{MMSE}}$, thereby completing the procedure for obtaining the MMSE estimate. A detailed proof is provided in Appendix D.
@@
We conclude the section by showing the next proposition. It shows that the MMSE and MAP estimators converge in the high SNR regime (i.e., $\sigma\to 0$), as demonstrated empirically in Figures 3, 4, and 5. This implies that the MMSE estimator’s superior statistical properties are most advantageous in low SNR conditions, which are common in structural biology applications like cryo-EM and cryo-ET Bendory et al. (2020). In these low SNR environments, the MMSE estimator consistently outperforms its MAP counterpart.
@@
The proof of the proposition is provided in Appendix E. The proof relies on the existing result, (Robert, 1999, Theorem 5.10), and is valid not only to rotations but also to other group operators, such as translations.
@@ 3 Numerical methods for MMSE orientation estimation
Since the posterior distribution of $g\in\mathsf{SO}(3)$ in (2.6) is continuous, a numerical discretization of the rotation group is required. In this work, the expectation is approximated using a quadrature rule over $\mathsf{SO}(3)$. Let $\{(g^{(\ell)},w_{\ell})\}_{\ell=0}^{L-1}$ denote a numerical quadrature on $\mathsf{SO}(3)$, constructed as a product of a spherical quadrature on $\mathsf{S}^{2}$ and a circular rule over the Euler angles Gräf and Potts (2009); Gräf (2012). The implementation employed here follows the publicly available MATLAB code described in Hoskins et al. (2024).
@@
Both estimators require evaluating $\|y-x_{\ell}\|$ for all candidate rotations, which naively costs $O(d)$ per candidate and thus $O(Ld)$ overall. This estimate, however, overlooks the structure of $\mathsf{SO}(3)$. Any 3D rotation can be decomposed into a viewing direction (two degrees of freedom, discretized with about $L^{2/3}$ samples) and an in-plane rotation about that direction (one degree of freedom, discretized with $L^{1/3}$ samples). By exploiting FFT-based methods, all $L^{1/3}$ in-plane rotations can be evaluated simultaneously for each viewing direction. Consequently, the effective complexity becomes $O(L^{2/3}\,d\,\log d)$, which improves over the naive bound by a factor of $L^{1/3}$ Kostelec and Rockmore (2008); Kileel et al. (2025a). This reduction is substantial in practice, since $L$ often reaches tens or even hundreds of thousands in high-resolution cryo-EM.
@@
One key advantage of the Bayesian framework is its flexibility in incorporating different prior distributions for the rotations. In practical cryo-EM applications, the distribution of particle orientations is often non-uniform due to preferred particle orientations or sample preparation artifacts Tan et al. (2017); Lyumkis (2019). If the rotation distribution can be well estimated in an early stage, the information of the rotation distribution can then be integrated into the Bayes estimator to improve rotation estimation.
To illustrate rotation estimation under different prior distributions on $\mathsf{SO}(3)$, we replace the uniform distribution on $\mathsf{SO}(3)$ with an isotropic Gaussian (IG) distribution on $\mathsf{SO}(3)$, denoted by $g\sim\mathcal{IG}_{\mathsf{SO}(3)}(\eta)$, parameterized by a scalar variance $\eta^{2}$. This distribution is frequently used in machine learning probabilistic models on $\mathsf{SO}(3)$ Corso et al. (2023); Leach et al. (2022); Jagvaral et al. (2024). It is worth noting that the IG distribution serves as a typical example, and similar phenomena as observed here extend to other non-uniform distributions on $\mathsf{SO}(3)$ as well.
The IG distribution $\mathcal{I}\mathcal{G}_{\mathsf{SO}(3)}(\eta)$ can be represented in an axis-angle form, with uniformly sampled axes of rotation and a rotation angle $\omega\in[0,\pi]$. The scalar variance $\eta^{2}$ controls the distribution of the rotation angle $\omega$: as $\eta\to\infty$, the $\mathsf{SO}(3)$ distribution approaches uniformity, whereas as $\eta\to 0$, $\omega$ becomes increasingly concentrated around 0, i.e., the rotation angle around the rotation axis is small. We apply the inverse sampling method to obtain i.i.d. samples from the IG distribution. Further details on this distribution are provided in Appendix F.1.
@@
We note that in experimental cryo-EM/ET datasets the viewing-angle distribution is typically unknown. Estimating this distribution, and understanding how prior misspecification impacts downstream reconstruction and heterogeneity analysis remains an important open problem. We discuss possible approaches for estimating the rotational prior in Appendix F.2. Recent work has begun to quantify misspecification effects under non-uniform group actions; see, e.g., Xu et al. (2025). In this light, Figure 2 can be viewed as a controlled prior-mismatch study: the data are generated from a concentrated isotropic-Gaussian prior, whereas the MMSE (and MAP) estimators are evaluated using an assumed prior with varying concentration. We observe a graceful degradation under mismatch, and a clear improvement as the assumed prior approaches the true distribution. Importantly, even under the common assumption of a uniform prior, the MMSE estimator improves over MLE in the low-SNR regime, while more accurate prior information yields additional gains when strong preferred orientations are present.
Figure 2: Impact of incorporating the prior rotation distribution on estimation accuracy. Simulations are performed for the cryo-ET model (1.2), excluding the projection step. The true rotation distribution is modeled as an isotropic Gaussian $\mathcal{IG}_{\mathsf{SO}(3)}(\eta=0.1)$. Estimation performance is measured using the geodesic distance defined in (2.8). Here, $g$ denotes the true rotation, $\hat{g}_{\mathrm{MLE}}$ is the maximum-likelihood estimator from (3.4), $\hat{g}_{\mathrm{MAP}}$ is the maximum a posteriori estimator from (2.11), and $\hat{g}_{\mathrm{MMSE}}$ denotes the Bayesian minimum mean square error estimator from (3.1). The MMSE estimators are computed assuming isotropic Gaussian priors on $\mathsf{SO}(3)$ with different concentration parameters $\eta\in\{0.5,0.1\}$ (see Appendix F.1). As $\eta$ decreases, the prior becomes more concentrated and closer to the true underlying distribution, leading to improved accuracy of both the MAP and MMSE estimators. Each data point in the plot is averaged over 3000 Monte Carlo trials using a rotation grid of size $L=2976$. Bottom $2\times 2$ panel. The four images compare denoised 3D volumes obtained using different rotation estimators at two representative noise regimes marked on the curve plot. Rows correspond to the estimator: top row uses $\hat{g}_{\mathrm{MLE}}$, bottom row uses $\hat{g}_{\mathrm{MMSE}}$ with the true prior $\mathcal{IG}_{\mathsf{SO}(3)}(\eta=0.1)$. Columns correspond to SNR: left column is a low-SNR example, and right column is a high-SNR example (where the true rotation is correctly classified).
@@
Sampling resolution and spectral information.
Another factor that influences the discrepancy between the MAP and MMSE estimators is the effective spectral information available at a given sampling resolution. When comparing MAP and MMSE estimates across different image sizes, it is important to distinguish increasing the number of pixels from increasing usable spectral content. If the underlying signal is effectively bandlimited and the image is already sampled at (or above) the Nyquist rate for that bandlimit, then increasing the sampling resolution mainly amounts to oversampling and should not materially change the likelihood or the posterior over $\mathsf{SO}(3)$; consequently, the MAP–MMSE estimators are expected to remain essentially unchanged. In contrast, when higher sampling resolution is accompanied by additional usable high-frequency signal (i.e., non-negligible per-frequency SNR at higher bands), the posterior typically becomes more concentrated, and we expect the MAP and MMSE estimates to approach each other in geodesic distance (at a fixed SNR level).
@@ 4 Bayesian orientation estimation as part of the volume reconstruction problem
Structure reconstruction typically follows two primary approaches: hard-assignment or soft-assignment methods, which are generally implemented through iterative refinement. In the hard-assignment approach, each observation is assigned a single orientation based on the highest correlation, and the 3D structure is then reconstructed given the rotations. In contrast, the soft-assignment method assigns probabilities across all possible orientations for each observation, enabling the 3D structure to be recovered as a weighted average of the observations, with weights determined by these probabilities. The iterative application of the soft-assignment procedure aligns with the EM algorithm Dempster et al. (1977); Sigworth et al. (2010), which serves as the core computational method in modern cryo-EM Scheres (2012b); Punjani et al. (2017).
@@ 4.1 Connection to the EM algorithm in volume reconstruction without projection
Unlike the previous orientation estimation model (2.1), we consider the following simplified model for volume reconstruction (see also Singer (2018); Bandeira et al. (2023); Fan et al. (2024)). We observe $M$ i.i.d samples taking the form
@@
where $V$ is 3D volume structure of interest, $\{g_{i}\}_{i=0}^{M-1}\in\mathsf{SO}(3)$ are unknown latent variables following i.i.d. uniform distribution, satisfying $g_{i}\circ V(x)\equiv V(g_{i}^{-1}x)$, and $\{\varepsilon_{i}\}_{i=0}^{M-1}$ are i.i.d isotropic Gaussian noise with variance $\sigma^{2}$. For the case where $V$ represents a 2D image, $\{g_{i}\}_{i=0}^{M-1}\in\mathsf{SO}(2)$ corresponds to in-plane rotations, and the model is used for image recovery (see more in Section 4.2Ma et al. (2019). The goal is to recover $V$ from the observations $\{y_{i}\}_{i=0}^{M-1}$, treating the rotations $\{g_{i}\}_{i=0}^{M-1}$ as latent variables.
@@
    $\hat{V}^{(t+1)}=$ $\underset{V}{\operatorname*{arg\,min}}\ \frac{1}{M}\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\left\|y_{i}-\left(g\circ V\right)\right\|^{2}\mathrm{d}g$
    $=$ $\frac{1}{M}\sum_{i=0}^{M-1}\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)(g^{-1}\circ y_{i})\mathrm{d}g.$ (4.3)
To state the next result, we first introduce the MMSE back-rotation operator. Given an observation $y_{i}$ and the current reference volume $\hat{V}^{(t)}$, let $p_{i}^{(t)}(g)$ denote the posterior weights over $g\in\mathsf{SO}(3)$ as defined in (4.2). We define $\widehat{\mathfrak{g}}_{\mathsf{MMSE},i,t}$ as the posterior average of the inverse action $g^{-1}$:
$\widehat{\mathfrak{g}}_{\mathsf{MMSE},i,t}:=\mathbb{E}\!\left[g^{-1}\,\big|\,y_{i},\hat{V}^{(t)}\right],$ (4.4)
which should be interpreted as a linear operator acting on functions (or volumes). Namely, for any test function $f$ on which rotations act,
$\big(\widehat{\mathfrak{g}}_{\mathsf{MMSE},i,t}\circ f\big):=\int_{g\in\mathsf{SO}(3)}p_{i}^{(t)}(g)\,\big(g^{-1}\circ f\big)\,\mathrm{d}g.$ (4.5)
Proposition 4.1 (MMSE operator form of the EM M-step).
Let $\hat{V}^{(t+1)}$ be the $(t+1)$-th volume estimator in the EM algorithm as described in Algorithm 2, and let $\widehat{\mathfrak{g}}_{\mathsf{MMSE},i,t}$ be defined as in (4.4)–(4.5). Then the M-step update can be written as
$\hat{V}^{(t+1)}=\frac{1}{M}\sum_{i=0}^{M-1}\widehat{\mathfrak{g}}_{\mathsf{MMSE},i,t}\circ y_{i}.$ (4.6)
The proof of this proposition is presented in Appendix G. In words, the proposition shows that the update rule for the volume structure estimation at iteration $t+1$, given the volume $\hat{V}^{(t)}$, is equivalent to aligning each observation $\left\{y_{i}\right\}_{i=0}^{M-1}$ using the associated back-rotation MMSE operator $\widehat{\mathfrak{g}}_{\text{MMSE},i,t}$, computed based on the reference volume $\hat{V}^{(t)}$, and then averaging the aligned observations (4.6). Thus, the MMSE back-rotation operator is a key ingredient in the EM algorithm for volume reconstruction.
@@
The MMSE back-rotation operator $\widehat{\mathfrak{g}}_{\mathsf{MMSE},i,t}$ (4.4) should be distinguished from the MMSE rotation matrix estimator $\hat{g}_{\mathsf{MMSE}}$ obtained in (2.12) by taking a posterior mean in the $3\times 3$ matrix embedding and then projecting onto $\mathsf{SO}(3)$ (via Orthogonal Procrustes). In the matrix formulation, one forms $\hat{g}_{\mathsf{relax}}$ as in (2.13) and then projects to obtain a single $\mathsf{SO}(3)$ element. In contrast, in (4.5) the inversion is handled inside the posterior average: the expectation is taken over the inverse actions $g^{-1}\circ(\cdot)$, yielding a linear operator that acts directly on $y_{i}$ (and is not, in general, itself a member of $\mathsf{SO}(3)$). While these two objects differ in representation, they are both posterior-weighted aggregations over rotations; the operator form is the one that naturally appears in the EM M-step for reconstructing the volume.
Despite the subtle differences arising from rotation estimation versus structure reconstruction, the previously introduced MMSE rotation estimator (2.12) can still serve as a practical tool for assessing the performance of soft assignments. Concretely, we take the posterior rotation matrix mean as in (4.4), and then perform an Orthogonal Procrustes projection to replace $\widehat{\mathfrak{g}}_{\mathsf{MMSE},i,t}$. In this way, the MMSE rotation estimator provides a concrete approximation to the soft assignment in structure reconstruction. In numerical experiments, it is compared with the hard assignment, demonstrating superior performance relative to the maximum-likelihood estimator, as discussed below.